# Visibility into AI Agents

### v5

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: The 2024 ACM Conference on Fairness, Accountability, and Transparency; June 3–6, 2024; Rio de Janeiro, Brazilbooktitle: The 2024 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’24), June 3–6, 2024, Rio de Janeiro, Brazildoi: 10.1145/3630106.3658948isbn: 979-8-4007-0450-5/24/06ccs: Computing methodologies Artificial intelligenceccs: Applied computing Law
@@ 1. Introduction
Many AI developers are creating systems with greater autonomy, access to external tools or services, and an increased ability to reliably adapt, plan, and act open-endedly over long time-horizons to achieve goals (Chan et al., 2023b; Richards, 2023; Ruan et al., 2023; Liu et al., 2023; Sharkey et al., 2023; Kenton et al., 2023; Sumers et al., 2023). We will say that such systems possess relatively high degrees of agency and will refer to them as (AI) agents or agentic systems (Lieberman, 1997; Chan et al., 2023b; Shavit et al., 2023). Systems with relatively low degrees of agency are those that only aid human decision-making or produce outputs without acting in the world, such as image classifiers or text-to-image models. Examples of agents could include reinforcement learning systems (Team et al., 2023a; Reed et al., 2022) that interact extensively with the real world [^1] Including the physical environment but also digital environments such as online platforms. or more capable versions of language models with tool or service access that could, for example, plan and book a holiday or send an email on a user’s behalf (OpenAI, 2023a; Schick et al., 2023; Ruan et al., 2023; Bran et al., 2023).
@@ 1.1.2. Overreliance and Disempowerment
Overreliance on AI agents to automate complex, high-stakes tasks could lead to severe consequences. Humans can already rely on certain automated systems more than is warranted (Cummings, 2004; Endsley and Kiris, 1995; Endsley and Kaber, 1999). More capable agents may enable automation of an increasing array of complex and useful tasks. Users—including both individuals and institutions—may rely on agents even in high-stakes situations, such as interfacing with the financial or legal systems, because human alternatives (e.g., hiring a lawyer) may become relatively slower and more expensive. At the same time, these agents may malfunction for a variety of reasons, including design flaws (Raji et al., 2022a; Ngo et al., 2022; Zwetsloot and Dafoe, 2019) or adversarial attack (Willison, 2023; Bailey et al., 2023; Zou et al., 2023). Malfunction may not be immediately apparent, especially if users lack the requisite expertise or domain knowledge. Stopping the agent may be difficult if doing so would lead to cascading failures or a competitive disadvantage for the user (Shavit et al., 2023). More broadly, profit and efficiency motives may lead to collective dependence on agents for essential societal functions, such as the provision of government services (De La Garza, 2020; Zilka et al., 2022) or the operation of essential infrastructure (Degrave et al., 2022; Biagioni et al., 2022). Companies providing access to AI agents would hold substantial power (Burrell and Fourcade, 2021), while malfunction of those agents could have societal-scale impacts.
At minimum, societies require information about the extent of reliance upon AI agents and whether such reliance is justified.
@@ 1.1.3. Delayed and Diffuse Impacts
Potential negative impacts of AI agents may be delayed and diffuse. [^4] Roughly speaking, we consider an impact to be diffuse if it is difficult to observe and most apparent in aggregate across many individual cases. Delayed and diffuse impacts may be difficult to manage because they may require sustained attention over long periods of time even to notice. Impacts of agents may be delayed if users give agents long-horizon goals, while diffuseness of impact may come from the widespread deployment of agents to automate complex processes. Consider an agent given the goal of continually finding and hiring job candidates who will most contribute to the company over the long-term. This agent may screen résumés (Gan et al., 2024), perform interviews, make the final hiring decision, and analyze the performance of hires. Given the time horizon over which the agent is acting and its influence over the company, any potential problems like algorithmic bias (Raghavan et al., 2020) could be hard to identify and become deeply entrenched. The most severe consequences of such problems may only be apparent when looking at how companies in aggregate use AI agents for hiring. AI agents could also subtly benefit their developers, akin to the self-preferencing behaviour of large-scale digital platforms (Kittaka et al., 2023). Moreover, agents that mediate or even substitute for human communication (noa, 2023b; Melley, 2024) could have diffuse and delayed psychological and social impacts (Jakesch et al., 2023; Bai et al., 2024; Karinshak et al., 2023), analogous to certain effects of social media (Bond et al., 2012; Sheffield and Francois, 2021; Milli et al., 2023). The deployment of agents may also induce changes in market structures or workforce impacts from job displacement (Acemoglu and Restrepo, 2018, 2019).
Identifying delayed and diffuse impacts may require long-term tracking of the extent and nature of AI agent usage across a wide range of application areas.
@@ 1.1.4. Multi-Agent Risks
Interactions and dependencies between many deployed agents could lead to risks not present at the level of a single system (Green, 2023; Park et al., 2023; Schelling, 1978; Hammond and TBD, 2024). Agents could enter into destabilising feedback loops, such as those between automated trading algorithms in the 2010 flash crash (CTFC and SEC, 2010). Agents partially built upon the same components—such as a particular foundation model—could have common vulnerabilities and failure modes (Bommasani et al., 2022; Cohen and Havlin, 2010); widespread deployment of such agents could risk large-scale systemic harms. More generally, there may be unpredictable behavioural changes that are characteristic of complex systems (Siegenfeld and Bar-Yam, 2020; Cohen and Havlin, 2010; Schelling, 1978). Competitive pressures and selection effects could lead to the development of agents that act in more anti-social ways (Xu et al., 2023; Hendrycks, 2023; Cassidy, 2009; Ely and Szentes, 2023).
These potential issues motivate understanding not just individual agents, but also interactions within groups of agents.
@@ 1.1.5. Sub-Agents
Agents could instantiate more agents to accomplish (components of) a task, which may magnify several of the risks discussed so far. It may be advantageous for an agent to create potentially specialized and more efficient sub-agents, especially if doing so is cheap and fast. For example, an agent could call copies of itself through an API, or itself train, fine-tune, or otherwise program another agent. Sub-agents could be problematic because they introduce additional points of failure; each sub-agent may itself malfunction, be vulnerable to attack, or otherwise operate in a way contrary to the user’s intentions. Stopping an agent from causing further harm might involve intervening not only on the agent, but also on any relevant sub-agents (Carey and Everitt, 2023; Soares et al., 2015). Yet, this process may be difficult because we lack methods for determining when an agent has created a sub-agent.
Information about the extent of sub-agent creation and operation can enable a better understanding of the significance of these risks.
@@ 2. Definitions
Agency is the degree to which an AI system acts directly in the world to achieve long-horizon goals, with little human intervention or specification of how to do so. An (AI) agent is a system with a relatively high degree of agency; we consider systems that mainly predict without acting in the world, such as image classifiers, to have relatively low degrees of agency. Examples of agents include reinforcement learning systems (Team et al., 2023a; Reed et al., 2022) that interact extensively with the real world [^6] Including the physical environment but also digital environments such as online platforms. or more capable versions of language models with tool access (OpenAI, 2023a; Schick et al., 2023; Ruan et al., 2023; Bran et al., 2023). We do not consider existing foundation models themselves to be agents. Our definitions compress the characterization of agency in Chan et al. (2023b), which points to four axes: the degree to which the system’s behaviour is specified, the degree to which the system’s behaviour is goal-directed, the degree to which the system has a direct impact in the world, and the degree to which the system can achieve goals over long time-horizons. For the purposes of this paper, we use agent and agentic system interchangeably.
@@
The (agent) deployer is the entity that operates an AI system and serves it to users. The deployer may not be the same as the developer(s). For example, Microsoft deploys OpenAI’s systems into its products (Warren, 2024), but did not develop GPT-4. A deployer may provide access to an agent in one of two ways. First, the deployer may serve a foundation model which users may combine with other components to make a more [^7] The system may not be completely autonomous since user approval may still be required for certain actions. agentic system. For example, users may use the scaffolding framework AutoGPT (Richards, 2023) to chain calls to GPT-4 (OpenAI, 2023b) and provide the model with tool access. Indeed, popular scaffolding frameworks depend upon an underlying foundation model (Wu et al., 2023; Richards, 2023). Second, the deployer may provide an agent or may furnish ways for users to make a provided system more agentic. For instance, OpenAI lets users build and share custom agents (OpenAI, 2024, 2023c) augmented with a variety of tools, including browsing, using Google Drive apps, and coding (OpenAI, 2023a).
@@
A tool or service refers to an external system or platform with which an AI agent interacts to perform its tasks. For instance, a flight booking website where the AI agent executes transactions, such as purchasing plane tickets on behalf of the user, would be a service. The tool provider or service provider is responsible for maintaining the platform that facilitates these interactions. We will often use tool and service interchangeably. Agents often interact with tools or services through dedicated APIs, which are interfaces and protocols specifically structured for agents, rather than human users.
@@ 3. Measures to Improve Visibility
We focus in this section on agents run by (agent) deployers—entities that deploy agents, or important subcomponents like a foundation model, as a service to users. We include foundation models because many frameworks for constructing agents use a foundation model as the central component (Wu et al., 2023; Richards, 2023; OpenAI, 2023a). While deployers are unlikely to account for all agent activity, they likely constitute a substantial fraction because the most capable foundation models are only available through deployers (OpenAI, 2023b; Team et al., 2023b; Bai et al., 2022), [^8] Currently, these deployers also happen to be developers. and using a deployer may be more convenient than running a system oneself. Moreover, since deployers can already see the inputs and outputs of deployed systems, they can attach agent identifiers to outputs, perform real-time monitoring, and collect activity logs. In Section 4, we analyze how to extend our measures to decentralized deployments of agents.
@@ 3.1.1. Types of Agent Identifiers
Which outputs contain the identifier? Decisions to attach an identifier may consider both the format and content of the output. By format, we mean whether the agent outputs data such as images, text, audio, or API requests to a service provider. An agent identifier’s specific implementation depends upon the output format. For example, identifiers for image outputs could be watermarks (Liu et al., 2024; Wang et al., 2021), while an identifier for an API request could be a simple header, similar to headers in HTTP requests. The difficulty of implementing identifiers varies based on the format of the output: for instance, adversarial users may easily remove watermarks (Zhang et al., 2023).
Regarding content, identification may be especially important for significant outputs. Certain outputs, such as purchases made on behalf of the user, may merit identifiers by virtue of the task the agent is accomplishing. Other outputs may only be significant beyond a certain threshold, such as requests for compute resources that exceed a certain amount.
@@ 3.1.2. Attaching Additional Information to Agent Identifiers
The specific instance of the agent. This information could consist of how the agent was deployed (e.g., by its user directly or by another agent); a list of external tools or services that the agent can access (e.g., applications or software, any physical tools); the scaffolding framework (e.g. memory or planning mechanisms); the intended scope, permissions, and goal(s) of the agent (Gilbert et al., 2023); or the sector of the intended deployment environment (e.g., finance). This type of information could be useful for regulators to develop an overall understanding of the extent and significance of agent use.
@@ 3.3.1. Agent-Specific Information
While activity logs are useful for understanding risks from AI systems in general (Javadi et al., 2020, 2021) and are common practice in many domains (ISO, 2022; noa, 2022, 2023a; Commission, 2022a, b), they can be especially helpful for addressing the risks we identified in Section 1.1. Different categories of logs could exist, such as for tool use, internal reasoning (Wei et al., 2023), self-critique (Saunders et al., 2022), memory management, etc, each of which may be useful to different actors.
@@ 4.1. Compute Provider Oversight
Compute providers could enable oversight over deployments using large quantities of compute. Such large-scale deployments could be concerning because they might involve vast numbers of agents, which could translate into a large impact multiplier for the user. Large-scale deployments are also noticeable because they consume significant resources. Users may be incentivized to employ compute providers for such deployments, rather than their own hardware, due to compute providers’ economies of scale; indeed, using compute as a service (e.g., infrastructure as a service or cloud) is the default way for a business to deploy its IT services. If a compute provider can identify large-scale deployments and whether they correspond to agent activities, they may ask the user for proof that they have implemented certain visibility measures (Egan and Heim, 2023; O’Brien et al., 2023).
@@ 4.3. Risks
Distributed enforcement mechanisms through tool and service providers also face similar concerns. If useful tools and services were closed off to agents that do not come from certified deployers, users may face strong pressure to use agents from such deployers. Whether because of government demand (Greenwald and MacAskill, 2013) or regulatory capture (Dal Bó, 2006), those deployers may have practises that are inimical to users or may not be responsive to user interests. Visibility measures for those deployers may be extremely invasive, such as excessive and unjustified logging. Agents from those deployers may not be well-suited to the user’s use cases; for example, the user might require an agent to be able to operate in a low-resource language. If the market of deployers is heavily concentrated, further reliance upon them could exacerbate systemic risks (Widder et al., 2023).
@@
Rather than mandating denial of requests from agents that do not have identifiers or that do not provide proof that certain visibility measures are implemented, one alternative is accounting for such compliance when determining the legal liability of users who deploy their own agents. Analogously, compliance with HIPAA de-identification standards can be taken into consideration to reduce regulatory fines or audits for violations (noa, ; of Human and Health Services, 2022). Similarly, the 2023 U.S. National Cybersecurity Strategy proposes to shield from private liability companies that follow cybersecurity best practices (House, 2023). Accounting for compliance when determining liability may incentivize standards adherence.

### v6

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: The 2024 ACM Conference on Fairness, Accountability, and Transparency; June 3–6, 2024; Rio de Janeiro, Brazilbooktitle: The 2024 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’24), June 3–6, 2024, Rio de Janeiro, Brazildoi: 10.1145/3630106.3658948isbn: 979-8-4007-0450-5/24/06ccs: Computing methodologies Artificial intelligenceccs: Applied computing Lawccs: Social and professional topics Governmental regulations
@@ 1. Introduction
Many AI developers are creating systems with greater autonomy, access to external tools or services, and an increased ability to reliably adapt, plan, and act open-endedly over long time-horizons to achieve goals (Chan et al., 2023b; Richards, 2023; Ruan et al., 2023; Liu et al., 2023; Sharkey et al., 2023; Kenton et al., 2023; Sumers et al., 2023). We will say that such systems possess relatively high degrees of agency and will refer to them as (AI) agents or agentic systems (Lieberman, 1997; Chan et al., 2023b; Shavit et al., 2023). Systems with relatively low degrees of agency are those that only aid human decision-making or produce outputs without acting in the world, such as image classifiers or text-to-image models. Examples of agents could include reinforcement learning systems (Team et al., 2023b; Reed et al., 2022) that interact extensively with the real world [^1] Including the physical environment but also digital environments such as online platforms. or more capable versions of language models with tool or service access that could, for example, plan and book a holiday or send an email on a user’s behalf (OpenAI, 2023a; Schick et al., 2023; Ruan et al., 2023; Bran et al., 2023).
@@ 1.1.2. Overreliance and Disempowerment
Overreliance on AI agents to automate complex, high-stakes tasks could lead to severe consequences. Humans can already rely on certain automated systems more than is warranted (Cummings, 2004; Endsley and Kiris, 1995; Endsley and Kaber, 1999). More capable agents may enable automation of an increasing array of complex and useful tasks. Users—including both individuals and institutions—may rely on agents even in high-stakes situations, such as interfacing with the financial or legal systems, because human alternatives (e.g., hiring a lawyer) may become relatively slower and more expensive. At the same time, these agents may malfunction for a variety of reasons, including design flaws (Raji et al., 2022a; Ngo et al., 2022; Zwetsloot and Dafoe, 2019) or adversarial attack (Willison, 2023; Bailey et al., 2023; Zou et al., 2023). Malfunction may not be immediately apparent, especially if users lack the requisite expertise or domain knowledge. Stopping the agent may be difficult if doing so would lead to cascading failures or a competitive disadvantage for the user (Shavit et al., 2023). More broadly, profit and efficiency motives may lead to collective dependence on agents for essential societal functions, such as the provision of government services (De La Garza, 2020; Zilka et al., 2022) or the operation of essential infrastructure (Degrave et al., 2022; Biagioni et al., 2022). Companies providing access to AI agents would hold substantial power (Burrell and Fourcade, 2021), while malfunction of those agents could have societal-scale impacts. At minimum, societies require information about the extent of reliance upon AI agents and whether such reliance is justified.
@@ 1.1.3. Delayed and Diffuse Impacts
Potential negative impacts of AI agents may be delayed and diffuse. [^4] Roughly speaking, we consider an impact to be diffuse if it is difficult to observe and most apparent in aggregate across many individual cases. Delayed and diffuse impacts may be difficult to manage because they may require sustained attention over long periods of time even to notice. Impacts of agents may be delayed if users give agents long-horizon goals, while diffuseness of impact may come from the widespread deployment of agents to automate complex processes. Consider an agent given the goal of continually finding and hiring job candidates who will most contribute to the company over the long-term. This agent may screen résumés (Gan et al., 2024), perform interviews, make the final hiring decision, and analyze the performance of hires. Given the time horizon over which the agent is acting and its influence over the company, any potential problems like algorithmic bias (Raghavan et al., 2020) could be hard to identify and become deeply entrenched. The most severe consequences of such problems may only be apparent when looking at how companies in aggregate use AI agents for hiring. AI agents could also subtly benefit their developers, akin to the self-preferencing behaviour of large-scale digital platforms (Kittaka et al., 2023). Moreover, agents that mediate or even substitute for human communication (noa, 2023b; Melley, 2024) could have diffuse and delayed psychological and social impacts (Jakesch et al., 2023; Bai et al., 2024; Karinshak et al., 2023), analogous to certain effects of social media (Bond et al., 2012; Sheffield and Francois, 2021; Milli et al., 2023). The deployment of agents may also induce changes in market structures or workforce impacts from job displacement (Acemoglu and Restrepo, 2018, 2019). Identifying delayed and diffuse impacts may require long-term tracking of the extent and nature of AI agent usage across a wide range of application areas.
@@ 1.1.4. Multi-Agent Risks
Interactions and dependencies between many deployed agents could lead to risks not present at the level of a single system (Green, 2023; Park et al., 2023; Schelling, 1978; Hammond and TBD, 2024). Agents could enter into destabilising feedback loops, such as those between automated trading algorithms in the 2010 flash crash (CTFC and SEC, 2010). Agents partially built upon the same components—such as a particular foundation model—could have common vulnerabilities and failure modes (Bommasani et al., 2022; Cohen and Havlin, 2010); widespread deployment of such agents could risk large-scale systemic harms. More generally, there may be unpredictable behavioural changes that are characteristic of complex systems (Siegenfeld and Bar-Yam, 2020; Cohen and Havlin, 2010; Schelling, 1978). Competitive pressures and selection effects could lead to the development of agents that act in more anti-social ways (Xu et al., 2023; Hendrycks, 2023; Cassidy, 2009; Ely and Szentes, 2023). These potential issues motivate understanding not just individual agents, but also interactions within groups of agents.
@@ 1.1.5. Sub-Agents
Agents could instantiate more agents to accomplish (components of) a task, which may magnify several of the risks discussed so far. It may be advantageous for an agent to create potentially specialized and more efficient sub-agents, especially if doing so is cheap and fast. For example, an agent could call copies of itself through an API, or itself train, fine-tune, or otherwise program another agent. Sub-agents could be problematic because they introduce additional points of failure; each sub-agent may itself malfunction, be vulnerable to attack, or otherwise operate in a way contrary to the user’s intentions. Stopping an agent from causing further harm might involve intervening not only on the agent, but also on any relevant sub-agents (Carey and Everitt, 2023; Soares et al., 2015). Yet, this process may be difficult because we lack methods for determining when an agent has created a sub-agent. Information about the extent of sub-agent creation and operation can enable a better understanding of the significance of these risks.
@@ 2. Definitions
Agency is the degree to which an AI system acts directly in the world to achieve long-horizon goals, with little human intervention or specification of how to do so. An (AI) agent is a system with a relatively high degree of agency; we consider systems that mainly predict without acting in the world, such as image classifiers, to have relatively low degrees of agency. Examples of agents include reinforcement learning systems (Team et al., 2023b; Reed et al., 2022) that interact extensively with the real world [^6] Including the physical environment but also digital environments such as online platforms. or more capable versions of language models with tool access (OpenAI, 2023a; Schick et al., 2023; Ruan et al., 2023; Bran et al., 2023). We do not consider existing foundation models themselves to be agents. Our definitions compress the characterization of agency in Chan et al. (2023b), which points to four axes: the degree to which the system’s behaviour is specified, the degree to which the system’s behaviour is goal-directed, the degree to which the system has a direct impact in the world, and the degree to which the system can achieve goals over long time-horizons. For the purposes of this paper, we use agent and agentic system interchangeably.
@@
The deployer is the entity that operates an AI system and serves it to users. The deployer may not be the same as the developer(s). For example, Microsoft deploys OpenAI’s systems into its products (Warren, 2024), but did not develop GPT-4. A deployer may provide access to an agent in one of two ways. First, the deployer may serve a foundation model which users may combine with other components to make a more [^7] The system may not be completely autonomous since user approval may still be required for certain actions. agentic system. For example, users may use the scaffolding framework AutoGPT (Richards, 2023) to chain calls to GPT-4 (OpenAI, 2023b) and provide the model with tool access. Indeed, popular scaffolding frameworks depend upon an underlying foundation model (Wu et al., 2023; Richards, 2023). Second, the deployer may provide an agent or may furnish ways for users to make a provided system more agentic. For instance, OpenAI lets users build and share custom agents (OpenAI, 2024, 2023c) augmented with a variety of tools, including browsing, using Google Drive apps, and coding (OpenAI, 2023a).
@@
A tool or service refers to an external system or platform with which an AI agent interacts to perform its tasks. For instance, a flight booking website where the AI agent executes transactions, such as purchasing plane tickets on behalf of the user, would be a service. The provider of the tool or service is responsible for maintaining the system or platform. We will often use tool and service interchangeably. Agents often interact with tools or services through dedicated APIs, which are interfaces and protocols specifically structured for agents, rather than human users.
@@ 3. Measures to Improve Visibility
We focus in this section on agents run by (agent) deployers—entities that deploy agents, or important subcomponents like a foundation model, as a service to users. We include foundation models because many frameworks for constructing agents use a foundation model as the central component (Wu et al., 2023; Richards, 2023; OpenAI, 2023a). While deployers are unlikely to account for all agent activity, they likely constitute a substantial fraction because the most capable foundation models are only available through deployers (OpenAI, 2023b; Team et al., 2023a; Bai et al., 2022), [^8] Currently, these deployers also happen to be developers. and using a deployer may be more convenient than running a system oneself. Moreover, since deployers can already see the inputs and outputs of deployed systems, they can attach agent identifiers to outputs, perform real-time monitoring, and collect activity logs. In Section 4, we analyze how to extend our measures to decentralized deployments of agents.
@@ 3.1.1. Types of Agent Identifiers
Which outputs contain the identifier? Decisions to attach an identifier may consider both the format and content of the output. By format, we mean whether the agent outputs data such as images, text, audio, or API requests to a service provider. An agent identifier’s specific implementation depends upon the output format. For example, identifiers for image outputs could be watermarks (Liu et al., 2024; Wang et al., 2021), while an identifier for an API request could be a simple header, similar to headers in HTTP requests. The difficulty of implementing identifiers varies based on the format of the output: for instance, adversarial users may easily remove watermarks (Zhang et al., 2023). Regarding content, identification may be especially important for significant outputs. Certain outputs, such as purchases made on behalf of the user, may merit identifiers by virtue of the task the agent is accomplishing. Other outputs may only be significant beyond a certain threshold, such as requests for compute resources that exceed a certain amount.
@@ 3.1.2. Attaching Additional Information to Agent Identifiers
The specific instance of the agent. This information could consist of how the agent was deployed (e.g., by its user directly or by another agent); a list of external tools or services that the agent can access (e.g., applications or software, any physical tools); the scaffolding framework (e.g. memory or planning mechanisms); the intended scope, permissions, and goal(s) of the agent (Gilbert et al., 2023); or the intended sector of deployment (e.g., finance). This type of information could be useful for regulators to develop an overall understanding of the extent and significance of agent use.
@@ 3.3.1. Agent-Specific Information
While activity logs are useful for understanding risks from AI systems in general (Javadi et al., 2020, 2021) and are common practice in many domains (ISO, 2022; noa, 2022, 2023a; Commission, 2022b, a), they can be especially helpful for addressing the risks we identified in Section 1.1. Different categories of logs could exist, such as for tool use, internal reasoning (Wei et al., 2023), self-critique (Saunders et al., 2022), memory management, etc, each of which may be useful to different actors.
@@ 4.1. Compute Provider Oversight
Compute providers could enable oversight over deployments that involve large quantities of compute. Large-scale deployments could be concerning because they might involve vast numbers of agents, which could translate into a large impact multiplier for the user. Large-scale deployments are also noticeable because they consume significant resources. Compute providers may have cost advantages over users deploying their own hardware because of economies of scale. Indeed, using compute as a service (e.g., infrastructure as a service or cloud) is the default way for a business to deploy its IT services. If a compute provider can identify large-scale deployments and whether they correspond to agent activities, they may ask the user for proof that they have implemented certain visibility measures (Egan and Heim, 2023; O’Brien et al., 2023).
@@ 4.3. Risks
Enforcement through tool and service providers also faces similar concerns. If useful tools and services were unavaiable to agents that were not from certified deployers, users may face strong pressure to use agents from such deployers. Whether because of government demand (Greenwald and MacAskill, 2013) or regulatory capture (Dal Bó, 2006), those deployers may have practises that are inimical to users or may not be responsive to user interests. Visibility measures for those deployers may be extremely invasive, such as excessive and unjustified logging. Agents from those deployers may not be well-suited to the user’s use cases; for example, the user might require an agent to be able to operate in a low-resource language. If the market of deployers is heavily concentrated, further reliance upon them could exacerbate systemic risks (Widder et al., 2023).
@@
Rather than mandating denial of requests from agents that do not have identifiers or that do not provide proof that certain visibility measures are implemented, one alternative is accounting for such compliance when determining the legal liability of users who deploy their own agents. Analogously, compliance with HIPAA de-identification standards can be taken into consideration to reduce regulatory fines or audits for violations (noa, [n. d.]; of Human and Health Services, 2022). Similarly, the 2023 U.S. National Cybersecurity Strategy proposes to shield from private liability companies that follow cybersecurity best practices (House, 2023). Accounting for compliance when determining liability may incentivize standards adherence.